OpenStudy (anonymous):

What is the solution to the equation –13 + x = –4? A. –17 B. 9 C. –9 D. 17

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Step 1: Simplify both sides of the equation. x−13=−4 Step 2: Add 13 to both sides. x−13+13=−4+13 x=9